OpenStudy (pradipgr817):

how can we find that a coin is biased just by tossing it a number of times.

OpenStudy (anonymous):

hey sunil this side i have problem in math. i want to solve this math if x+1/x=3 then the value of (x3+1/x)/x2-x+1 is =? solve this math

OpenStudy (anonymous):

Please read here : http://en.wikipedia.org/wiki/Checking_whether_a_coin_is_fair

OpenStudy (anonymous):

hey sunil this side i have problem in math. i want to solve this math if x+1/x=3 then the value of (x3+1/x)/x2-x+1 is =? solve this math